# On the Robustness of Kolmogorov-Arnold Networks: An Adversarial Perspective

### v2

@@ 4.2 Convolutional Models
In the context of iterative white-box adversarial attacks, CKANs and CNNs exhibit contrasting behaviors across datasets. As observed in Table 6, both CKANs and CNNs exhibit poor performance on the complex datases (CIFAR-10 and SVHN), demonstrating significant vulnerability to adversarial perturbations. On CIFAR-10, none of the models successfully classify any images generated by those attacks. On the SVHN dataset, the medium and large models manage to classify a small number of adversarially generated images. This suggests that as the dataset complexity increases, the models become more susceptible to adversarial attacks, likely due to the intricate decision boundaries required for accurate classification.
Table 6: White-box and black-box attacks on convolutional models, evaluated across MNIST, CIFAR-10, and SVHN datasets. The table reports robust accuracy for each attack, while the ‘Acc‘ column indicates the model’s accuracy on clean, unperturbed inputs. Boldface values indicate the most robust model in each experiment.
Dataset Model White-Box Black-Box Acc
FGSM PGD C&W MIM Square SimBA NES
MNIST $\text{CKAN}_{\text{small}}$ 72.36 4.52 5.70 8.98 38.39 34.82 49.90 99.47
$\text{CNN}_{\text{small}}$ 68.20 0.74 0.98 2.26 27.38 9.82 33.73 99.23
$\text{CKAN}_{\text{medium}}$ 86.33 29.88 33.12 42.72 57.93 41.76 66.07 99.55
$\text{CNN}_{\text{medium}}$ 28.36 0.00 0.00 0.00 0.39 7.24 8.92 99.36
$\text{CKAN}_{\text{large}}$ 84.72 31.28 34.32 39.64 56.84 36.21 58.92 99.52
$\text{CNN}_{\text{large}}$ 55.87 0.01 0.13 0.19 5.15 2.38 5.45 99.49
CIFAR-10 $\text{CKAN}_{\text{small}}$ 0.01 0.00 0.00 0.00 0.00 2.48 0.99 74.70
$\text{CNN}_{\text{small}}$ 0.01 0.00 0.00 0.00 0.00 2.28 0.69 73.03
$\text{CKAN}_{\text{medium}}$ 0.57 0.00 0.00 0.00 0.39 4.26 1.88 80.72
$\text{CNN}_{\text{medium}}$ 0.27 0.00 0.00 0.00 0.09 2.28 0.79 76.43
$\text{CKAN}_{\text{large}}$ 0.85 0.00 0.00 0.00 0.19 4.06 1.98 82.82
$\text{CNN}_{\text{large}}$ 1.94 0.00 0.00 0.00 1.48 4.46 2.38 79.26
SVHN $\text{CKAN}_{\text{small}}$ 1.04 0.00 0.00 0.00 7.44 10.41 7.44 92.34
$\text{CNN}_{\text{small}}$ 0.43 0.00 0.00 0.00 4.76 5.15 2.87 90.69
$\text{CKAN}_{\text{medium}}$ 6.74 0.02 0.02 0.02 16.96 16.36 13.59 94.37
$\text{CNN}_{\text{medium}}$ 5.00 0.05 0.05 0.04 14.78 9.52 8.03 93.41
$\text{CKAN}_{\text{large}}$ 8.33 0.09 0.10 0.09 19.84 16.26 13.09 94.41
$\text{CNN}_{\text{large}}$ 6.67 0.04 0.05 0.04 14.08 8.23 7.83 92.72
@@
This disparity is further corroborated by Figure 3, which reveals a faster loss convergence on complex datasets (CIFAR-10 and SVHN) compared to MNIST during the attack process. The rapid convergence of the loss implies that adversarial attacks “fool” models more easily on these complex datasets.
Figure 3: Comparative loss dynamics of CNNs and CKANs under PGD, MIM, and C&W attacks on the MNIST, CIFAR-10, and SVHM datasets. Each line represents the mean loss across batches per iteration, with shaded areas indicating the standard deviation.
For the non-iterative white-box attack (FGSM), the results are more definitive. On CIFAR-10, CKANs exhibit greater robustness than CNNs for medium-sized models, whereas CNNs perform better for large models. On the SVHN dataset, CKANs demonstrate slightly higher robustness across all model sizes, although the gap between CKANs and CNNs is relatively small. On the MNIST dataset, which is less complex, CKANs show significantly higher robustness than CNNs, with a noticeable gap favoring CKANs. As shown in Figure 4, this gap is not consistently large across different $\epsilon$ values.
@@
When analyzing black-box attacks, CKANs consistently demonstrated superior robustness compared to CNNs across most datasets and model sizes. CKANs outperformed CNNs in nearly all scenarios, except for the black-box attacks on large models for the CIFAR-10 dataset, where CNNs exhibited slightly better robustness, and on the small models for the CIFAR-10 dataset, where both of the configurations failed to classify any adversarial image generated with the Square attack.
@@
Table 7: Maximum transferability of our convolutional models for MNIST, CIFAR-10, and SVHM. The value in row $i$, column $j$ represents the maximum transferability between row-$i$ model and column-$j$ model across four attacks—FGSM, PGD, C&W, MIM—calculated using Equation 11. Unlike robustness tables, the values here indicate the percentage of images misclassified. The ‘Average’ row shows the average transferability for the attacking model. Boldface values denoting the lowest transferability, where a lower value indicates better robustness.
@@
$\text{CKAN}_{\text{small}}$ 95.21 2.94 3.26 4.70 20.6 4.80
$\text{CKAN}_{\text{medium}}$ 5.30 74.81 17.78 3.32 18.95 12.01
$\text{CKAN}_{\text{large}}$ 3.75 14.95 72.80 2.51 15.16 9.93
$\text{CNN}_{\text{small}}$ 3.43 1.71 1.55 99.22 33.47 6.72
$\text{CNN}_{\text{medium}}$ 3.82 2.78 3.48 7.80 100.00 46.28
$\text{CNN}_{\text{large}}$ 3.35 4.49 5.28 5.64 6.98 100.00
Average ($\downarrow$) 19.14 16.94 17.35 20.53 32.52 29.95
@@
$\text{CKAN}_{\text{small}}$ 100.00 53.39 45.62 62.95 60.19 35.45
$\text{CKAN}_{\text{medium}}$ 70.25 100.00 93.48 79.12 90.40 68.52
$\text{CKAN}_{\text{large}}$ 62.80 92.13 100.00 75.85 89.65 73.39
$\text{CNN}_{\text{small}}$ 70.47 71.27 69.74 100.00 85.70 59.78
$\text{CNN}_{\text{medium}}$ 63.23 82.31 82.52 82.50 100.00 78.68
$\text{CNN}_{\text{large}}$ 47.53 68.93 76.86 65.07 84.74 100.00
Average ($\downarrow$) 69.04 78.00 78.03 77.58 85.15 69.30
@@
$\text{CKAN}_{\text{small}}$ 100.00 63.08 58.62 72.23 58.73 57.70
$\text{CKAN}_{\text{medium}}$ 77.44 100.00 83.76 69.97 75.40 75.40
$\text{CKAN}_{\text{large}}$ 76.54 86.43 99.90 69.42 76.23 75.67
$\text{CNN}_{\text{small}}$ 68.52 55.79 52.85 100.00 63.33 62.15
$\text{CNN}_{\text{medium}}$ 72.11 74.21 71.98 78.26 99.96 85.87
$\text{CNN}_{\text{large}}$ 70.93 74.52 72.70 78.62 85.89 99.96
Average ($\downarrow$) 77.59 75.67 73.30 78.08 76.58 76.12
@@ 5 Concluding Remarks
The introduction of KANs opens new avenues in adversarial machine learning, presenting both challenges and opportunities for improving model security. In this study, we explored several key aspects of KANs, focusing on their behavior in convolutional and fully connected layers, which have not been extensively investigated before.
@@
Despite these promising findings, it is crucial to recognize that KANs remain vulnerable to adversarial perturbations. Our analysis reveals that while larger configurations enhance robustness, KANs are not impervious to adversarial attacks. This ongoing fragility underscores the need for continued research to develop more effective strategies for improving the security of KANs. Addressing these vulnerabilities is essential for the reliable deployment of KANs in security-critical applications.
As the field continues to advance, further investigation is required to better understand the strengths and limitations of FCKANs and CKANs. This research will be instrumental in paving the way for the practical application of KANs in domains where security and reliability are paramount.
@@ Appendix A Image Samples of Attacks
Convolutional Models. Figure 7 and Figure 8 illustrate adversarial images and perturbations for CNNs and CKANs under white-box and black-box attacks, showcasing architectural influences on adversarial patterns.

### v3

@@ 4.2 Convolutional Models
In the context of iterative white-box adversarial attacks, CKANs and CNNs exhibit contrasting behaviors across datasets. As observed in Table 7, both CKANs and CNNs exhibit poor performance on the multi-channel datasets (CIFAR-10, SVHN and ImageNet), demonstrating significant vulnerability to adversarial perturbations. On CIFAR-10, none of the models successfully classify any images generated by those attacks. On the SVHN dataset, the medium and large models manage to classify a small number of adversarially generated images. For ImageNet, the CKAN models failed to classify any images, exhibiting the same behavior observed on CIFAR-10. However, CNNs on medium and large scales did not fail completely, showing some level of robustness compared to their CKAN coutnerparts.
Table 7: White-box, black-box, and ensemble attack on convolutional models, evaluated across MNIST, CIFAR-10, SVHN, and ImageNet datasets. The table reports robust accuracy for each attack, while the Acc column indicates the model’s accuracy on clean, unperturbed inputs. Boldface values indicate the most robust model in each experiment. (For ImageNet NES attacks, denoted by *, results could not be fully evaluated as in Ilyas et al. (2018) due to computational constraints at our university.)
Dataset Model White-Box Black-Box Ensemble-Based Acc
FGSM PGD C&W MIM Square SimBA NES Auto Attack
MNIST $\text{CKAN}_{\text{small}}$ 72.36 4.52 5.70 8.98 38.39 34.82 49.90 1.13 99.47
$\text{CNN}_{\text{small}}$ 68.20 0.74 0.98 2.26 27.38 9.82 33.73 0.02 99.23
$\text{CKAN}_{\text{medium}}$ 86.33 29.88 33.12 42.72 57.93 41.76 66.07 14.78 99.55
$\text{CNN}_{\text{medium}}$ 28.36 0.00 0.00 0.00 0.39 7.24 8.92 0.00 99.36
$\text{CKAN}_{\text{large}}$ 84.72 31.28 34.32 39.64 56.84 36.21 58.92 18.89 99.52
$\text{CNN}_{\text{large}}$ 55.87 0.01 0.13 0.19 5.15 2.38 5.45 0.00 99.49
CIFAR-10 $\text{CKAN}_{\text{small}}$ 0.01 0.00 0.00 0.00 0.00 2.48 0.99 0.00 74.70
$\text{CNN}_{\text{small}}$ 0.01 0.00 0.00 0.00 0.00 2.28 0.69 0.00 73.03
$\text{CKAN}_{\text{medium}}$ 0.57 0.00 0.00 0.00 0.39 4.26 1.88 0.00 80.72
$\text{CNN}_{\text{medium}}$ 0.27 0.00 0.00 0.00 0.09 2.28 0.79 0.00 76.43
$\text{CKAN}_{\text{large}}$ 0.85 0.00 0.00 0.00 0.19 4.06 1.98 0.00 82.82
$\text{CNN}_{\text{large}}$ 1.94 0.00 0.00 0.00 1.48 4.46 2.38 0.00 79.26
SVHN $\text{CKAN}_{\text{small}}$ 1.04 0.00 0.00 0.00 7.44 10.41 7.44 0.00 92.34
$\text{CNN}_{\text{small}}$ 0.43 0.00 0.00 0.00 4.76 5.15 2.87 0.00 90.69
$\text{CKAN}_{\text{medium}}$ 6.74 0.02 0.02 0.02 16.96 16.36 13.59 0.01 94.37
$\text{CNN}_{\text{medium}}$ 5.00 0.05 0.05 0.04 14.78 9.52 8.03 0.03 93.41
$\text{CKAN}_{\text{large}}$ 8.33 0.09 0.10 0.09 19.84 16.26 13.09 0.06 94.41
$\text{CNN}_{\text{large}}$ 6.67 0.04 0.05 0.04 14.08 8.23 7.83 0.02 92.72
ImageNet $\text{CKAN}_{\text{small}}$ 0.20 0.00 0.00 0.00 59.40 70.60 $76.20^{*}$ 0.00 76.20
$\text{CNN}_{\text{small}}$ 1.60 0.00 0.00 0.00 58.80 74.60 $79.80^{*}$ 0.00 79.80
$\text{CKAN}_{\text{medium}}$ 4.00 0.00 0.00 0.00 45.00 64.60 $78.20^{*}$ 0.00 78.20
$\text{CNN}_{\text{medium}}$ 3.00 0.20 0.20 0.20 41.80 63.60 $77.40^{*}$ 0.20 77.40
$\text{CKAN}_{\text{large}}$ 6.40 0.00 0.00 0.00 22.20 37.20 $72.80^{*}$ 0.80 72.80
$\text{CNN}_{\text{large}}$ 11.20 0.80 1.20 1.00 34.80 47.00 $73.40^{*}$ 0.60 73.40
@@
This disparity is further corroborated by Figure 3, which reveals a faster loss convergence on complex datasets (CIFAR-10, SVHN and ImageNet) compared to MNIST during the attack process. The rapid convergence of the loss implies that adversarial attacks “fool” models more easily on these complex datasets.
Figure 3: Comparative loss dynamics of CNNs and CKANs under PGD, MIM, and C&W attacks on the MNIST, CIFAR-10, SVHM and ImageNet datasets. Each line represents the mean loss across batches, per iteration, with shaded areas indicating the standard deviation. For the MNIST, CIFAR-10 AND SVHN datasets, the shaded areas represent the original standard deviation, whereas for ImageNet, they represent the logarithm of the original std
For the non-iterative white-box attack (FGSM), the results are more definitive. On CIFAR-10, CKANs exhibit greater robustness than CNNs for medium-sized models, whereas CNNs perform better for large models. On ImageNet, small and large CNNs achieve greater robustness than CKANs. However, similar to CIFAR-10, the medium CKAN model demonstrates greater robustness than its CNN counterpart. On the SVHN dataset, CKANs demonstrate slightly higher robustness across all model sizes, although the gap between CKANs and CNNs is relatively small. On the MNIST dataset, which is less complex, CKANs show significantly higher robustness than CNNs, with a noticeable gap favoring CKANs. As shown in Figure 4, this gap is not consistently large across different $\epsilon$ values.
@@
When analyzing black-box attacks, for MNIST, CIFAR-10, and SVHN, CKANs consistently demonstrated superior robustness compared to CNNs across most datasets and model sizes. CKANs outperformed CNNs in nearly all scenarios, except for the black-box attacks on large models for the CIFAR-10 dataset, where CNNs exhibited slightly better robustness, and on the small models for the CIFAR-10 dataset, where both of the configurations failed to classify any adversarial image generated with the Square attack. For ImageNet, all models performed well under black-box attacks. However, for large models, CNNs significantly outperformed CKANs in terms of robustness.
In the case of AutoAttack on the MNIST dataset, CKANs performed exceptionally well against that attack. This suggests that fooling a large CKAN model is particularly challenging when dealing with grayscale images, potentially due to the unique structural properties of CKANs, which enhance their resistance to adversarial perturbations in such domains. In contrast, AutoAttack successfully “fooled” most of the models in the other datasets. On CIFAR-10, none of the models successfully classified even a single adversarial image generated by this attack. On SVHN and ImageNet, the medium CNN outperformed the medium CKAN, while the large CKAN performed better than the large CNN. The small models, similar to the CIFAR-10 models, failed to classify even a single image.
@@
Table 8: Maximum transferability of our convolutional models for MNIST, CIFAR-10, and SVHM. Each row represents the model that generates the adversarial example, while each column represents the model that evaluates those examples. The value in row $i$, column $j$ represents the maximum transferability between row-$i$ model and column-$j$ model across four attacks—FGSM, PGD, C&W, MIM—calculated using Equation 12. Unlike robustness tables, the values here indicate the percentage of images misclassified. The ‘Average’ row shows the average transferability for the attacking model. Boldface values denoting the lowest transferability, where a lower value indicates better robustness.
@@
$\text{CKAN}_{\text{small}}$ 2.94 3.26 4.70 20.6 4.80
$\text{CKAN}_{\text{medium}}$ 5.30 - 17.78 3.32 18.95 12.01
$\text{CKAN}_{\text{large}}$ 3.75 14.95 - 2.51 15.16 9.93
$\text{CNN}_{\text{small}}$ 3.43 1.71 1.55 - 33.47 6.72
$\text{CNN}_{\text{medium}}$ 3.82 2.78 3.48 7.80 - 46.28
$\text{CNN}_{\text{large}}$ 3.35 4.49 5.28 5.64 6.98 -
Average ($\downarrow$) 3.93 5.37 6.27 4.79 19.03 15.94
@@
$\text{CKAN}_{\text{small}}$ 53.39 45.62 62.95 60.19 35.45
$\text{CKAN}_{\text{medium}}$ 70.25 - 93.48 79.12 90.40 68.52
$\text{CKAN}_{\text{large}}$ 62.80 92.13 - 75.85 89.65 73.39
$\text{CNN}_{\text{small}}$ 70.47 71.27 69.74 - 85.70 59.78
$\text{CNN}_{\text{medium}}$ 63.23 82.31 82.52 82.50 - 78.68
$\text{CNN}_{\text{large}}$ 47.53 68.93 76.86 65.07 84.74 -
Average ($\downarrow$) 62.85 73.60 73.64 73.09 82.13 63.16
@@
$\text{CKAN}_{\text{small}}$ 63.08 58.62 72.23 58.73 57.70
$\text{CKAN}_{\text{medium}}$ 77.44 - 83.76 69.97 75.40 75.40
$\text{CKAN}_{\text{large}}$ 76.54 86.43 - 69.42 76.23 75.67
$\text{CNN}_{\text{small}}$ 68.52 55.79 52.85 - 63.33 62.15
$\text{CNN}_{\text{medium}}$ 72.11 74.21 71.98 78.26 - 85.87
$\text{CNN}_{\text{large}}$ 70.93 74.52 72.70 78.62 85.89 -
Average ($\downarrow$) 73.10 70.80 67.98 73.70 71.91 71.35
Our experiments indicate that smaller KANs do not consistently outperform conventional neural networks in adversarial settings. By contrast, once the KAN architecture is scaled, the decomposition advantage becomes more pronounced. These findings cohere with the theoretical conjecture stated in Section 2.3.1, namely, that having sufficient capacity allows KANs to learn smoother local transformations, distributing the learned decision boundary in a way that reduces susceptibility to adversarial perturbations.
While we do not claim a fully rigorous theory of adversarial robustness for KANs, these lines of reasoning provide plausible explanations for our empirical results, motivating further research on KAN-based Lipschitz regularization and other theoretical investigations.
4.3 Ablation Study: Impact of KAN Hyperparameters
To evaluate the impact of KAN hyperparameters on adversarial robustness, we analyzed the five configurations tested in our ablation study (Section 3.3).
The results in Figure 5 reveal an interesting phenomenon. In small models, across all datasets, the model with the highest spline order consistently achieves better robustness against adversarial attacks, except for FGSM, where a model with a spline order of 5 performs best. This trend remains stable in medium-sized models, where we observe that the model with the lowest spline order (3) performs significantly worse at correctly predicting adversarial examples.
Figure 5: Robust accuracy of FCKANs under different adversarial attacks, varying the number of knots and spline order.
However, when moving to large models, this trend shifts. In this case, models with the lowest spline order demonstrate strong adversarial robustness. Notably, for iterative white-box attacks, the model with 7 knots and a spline order of 3 exhibits remarkable stability, showing strong resistance to these attacks.
4.4 Adversarial training
Examining the results in Table 9 for the fully connected models, and Table 10 for the convolutional models, we can clearly see that, although the standard FCNNs and CNNs achieved higher accuracy on clean test-set images (except for the large models), they perform worse than their KAN-bases counterparts (FCKANs and FCKANs) when predicting adversarially perturbed images. This trend holds across all attack types, indicating that FCKAN and CKAN models are inherently more robust to adversarial perturbations compared to traditional neural networks. Furthermore, this improved robustness may stem from the ability of FCKANs and CKANs to better adapt to adversarial perturbations after exposure to adversarial examples during training. Their structured functional representations might enable them to generalize more effectively to unseen adversarial examples compared to FCNNs and CNNs, which struggle to adjust beyond the specific perturbations encountered during training.
Table 9: Adversarial attack results on fully connected models trained on KMNIST. Values represent robust accuracy (%) after each attack. The highest robust accuracy per row is boldfaced.
Model FGSM PGD C&W MIM Square SimBA NES Clean Acc
$\text{FCKAN}_{\text{small}}$ 68.5 62.14 60.87 62.41 55.75 54.16 60.61 79.24
$\text{FCNN}_{\text{small}}$ 54.63 49.12 47.78 49.71 52.97 47.42 57.73 85.56
$\text{FCKAN}_{\text{medium}}$ 77.12 72.38 71.73 72.85 71.03 69.44 74.20 88.95
$\text{FCNN}_{\text{medium}}$ 72.73 65.95 66.60 67.33 70.63 63.69 73.01 95.22
$\text{FCKAN}_{\text{large}}$ 64.72 57.82 57.21 58.55 53.17 51.88 56.54 74.87
$\text{FCNN}_{\text{large}}$ 48.05 41.92 41.7 43.21 46.42 40.87 50.09 74.08
Table 10: Adversarial attack results on convolutional models trained on CIFAR-10. Values represent robust accuracy (%) after each attack. The highest robust accuracy per row is boldfaced.
Model FGSM PGD C&W MIM Square SimBA NES Clean Acc
$\text{CKAN}_{\text{small}}$ 35.81 31.29 29.64 31.48 50.19 35.61 51.78 64.31
$\text{CNN}_{\text{small}}$ 30.28 23.68 23.80 23.96 46.62 26.88 46.82 65.97
$\text{CKAN}_{\text{medium}}$ 39.58 34.61 33.44 34.96 52.28 39.58 54.06 65.80
$\text{CNN}_{\text{medium}}$ 28.55 21.48 21.61 21.55 44.24 27.18 44.14 66.27
$\text{CKAN}_{\text{large}}$ 40.56 35.56 34.77 35.73 52.38 40.57 51.88 66.60
$\text{CNN}_{\text{large}}$ 31.74 25.02 24.73 25.12 45.23 28.57 47.22 66.42
@@ 5 Concluding Remarks
The introduction of KANs opens new avenues in adversarial machine learning, presenting both challenges and opportunities for improving model safety. In this study, we explored several key aspects of KANs, focusing on their behavior in convolutional and fully connected layers, which have not been extensively investigated before.
@@
Our ablation study further highlights the impact of key hyperparameters, specifically the number of knots and spline order, on adversarial robustness. We found that models with a higher spline order tend to be more resilient to adversarial perturbations, especially in smaller and medium-sized configurations. However, for large-scale models, lower spline order configurations demonstrated stronger robustness under iterative white-box attacks, suggesting that the interplay between network complexity and attack resilience is non-trivial. These findings indicate that optimizing KAN hyperparameters could be a promising avenue for enhancing robustness.
Moreover, our adversarial training experiments demonstrate that FCKANs and CKANs benefit significantly from exposure to adversarial examples during training. Notably, adversarially trained KAN models consistently outperformed their standard counterparts in robustness across all tested attacks. However, even after adversarial training, the performance gap between KANs and conventional architectures persisted in some settings, suggesting that additional regularization techniques or training strategies could further enhance their resilience.
While these results provide valuable insights, they also raise important open questions. Future work should explore whether alternative training methodologies—such as certified defenses, Lipschitz regularization, or hybrid adversarial training—can further improve KAN robustness.
Despite their theoretical advantages, KANs are not inherently immune to adversarial perturbations. However, their structured functional decomposition offers a unique framework that may be leveraged for more robust and interpretable deep learning models. As adversarial machine learning research advances, continued investigation into the interplay between KAN structure, optimization, and adversarial robustness will be crucial for their practical adoption in security-critical applications.
@@ Appendix A Image Samples of Attacks
Convolutional Models. Figure 8 and Figure 9 illustrate adversarial images and perturbations for CNNs and CKANs under white-box and black-box attacks, showcasing architectural influences on adversarial patterns. Similarly, Figure 10 illustrates adversarial images and perturbations for the ImageNet dataset under white-box attacks, showcasing the impact of different architectures on adversarial behavior.
@@
Figure 10: Visualization of the adversarial images and the perturbations generated by white-box attacks on three images from our subset of ImageNet using convolutional models. Rows represent model types and sizes: $\text{CNN}_{\text{Large}}$, $\text{CKAN}_{\text{Large}}$, $\text{CNN}_{\text{Medium}}$, $\text{CKAN}_{\text{Medium}}$, $\text{CNN}_{\text{Small}}$, and $\text{CKAN}_{\text{Small}}$. The first four images in each row represent the adversarial images, while the last four images represent the perturbations.